A persons lands $10000 to a corporation by purchasing a bond from the corporation. Simple interest is computed quarterly at a rate of 3 percent per quarter, and a check for the interest is mailed each quarter to all bondholders. The bond expires at the end of 5 years, and the final check includes the original principal plus interest earned during the last quarter. Compute the interest earned each quarter and the total interest will be earned over the 5 year life of the bond.
1
Expert's answer
2020-08-19T19:07:38-0400

Interest earned each quarter: "i=10000*0.03=\$300."

Total interest earned over 5 years: "I=i*5*4=300*5*4=\$6000."